OpenStudy (anonymous):

heres my problem..i have the answer and the example shows to put it in square root form but then says round to the nearest 100th. that's where im lost. A 15ft ladder is places against a verticle wall of a building with the bottom of the ladder standing level ground 14ft from the base of the building. how high up the wall does the ladder reach. I got 29. how do I do the square root of 29 rounded to the nearest 100th???

OpenStudy (amistre64):

by hand or by calculator ....

OpenStudy (anonymous):

is my answer 29 right? I tried and do it with a program I have on my computer but it only gives me the answer square root 29

OpenStudy (amistre64):

the set up for (x+y)^2 is x^2 +2xy + y^2, so lets find an x^2 to start with ---------- /29 . 00 00 5^2 = 25, 6^2 = 36 ... lets use 5 5 ---------- /29 . 00 00 25 ----- 4 00 now all we need to determine is 2xy + y^2, that equals 400, x=5 sooo, lets double the 5 and take on a y 5 ---------- /29 . 00 00 25 ----- 10_/ 4 00 , 4 is to big, lets try 3 5 . 3 ---------- /29 . 00 00 25 ----- 103/ 4 00 3 09 ------ 91 00 double the new x and find a y 5 . 3 ---------- /29 . 00 00 25 ----- 103/ 4 00 3 09 ------ 109_/ 91 00 , looks like 8 might work 5 . 3 8 ---------- /29 . 00 00 25 ----- 103/ 4 00 3 09 ------ 1098/ 91 00 87 84 ------ 3 16 ... just need one more to rnd eh

OpenStudy (amistre64):

by calculator we get sqrt(29)= 5.38516480713

OpenStudy (anonymous):

how did you do it by calculator?

OpenStudy (amistre64):

depends on the calculator .... you need one that can do exponents, or has a sqrt button

OpenStudy (anonymous):

ok what would 27 be?

OpenStudy (anonymous):

this e gets on my nerves. its sooo slow. ok I sent a message as well it shows the answer as 5.19615 so rounding to nearest 100th would that be 5.29